lolokrat:

Find the coordinates of the other endpoint of the segment with the given endpoint and midpoint. Endpoint: (0,11) Midpoint: (9,10)

Shadow:

these brothers trying to make our head work

lolokrat:

i have the formula (x1+x2)/2 , (y1+y2)/2 written on my notes

Shadow:

the way I see it, from the endpoint, it rose by -1 and ran by +9 so we do that again and we get our other end point

Shadow:

so -1 on the y axis and +9 on the x axis

lolokrat:

how do i plug that in the formula my teacher whack and wants work shown

Shadow:

Oh that's the formula for the midpoint but I just went through a lecture so I told you how to do it the dumb way lol

Shadow:

umm so you gotta separate the y and x coordinates. \[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\] \[x = \frac{ x _{1} + x _{2}}{ 2 }\] \[y = \frac{ y _{1} + y _{2} }{ 2 }\]

Shadow:

Do you see what you have to do?

lolokrat:

yea i wrote tht down like tht already

lolokrat:

idk how to work it out

Shadow:

what is 'x' in this case

Shadow:

and 'y'

lolokrat:

what

Shadow:

think about what this formula is for

Shadow:

Look at my bottom two equations

lolokrat:

x=x1+x2/2 x=0+x2/2 y=y1+y2/2 y=11+y2/2

lolokrat:

i have an example from class too

Shadow:

Yeah so you inputted the first endpoint but what are you solving for

Shadow:

just trust me with this

lolokrat:

the other endpoint

Shadow:

and you can't solve for the other endpoint if you have two unknown variables (x and x2). x2 is the x coordinate of the other endpoint, so what is x?

Shadow:

in (x1 + x2)/2 = x

lolokrat:

what x=x1+x2/2 x=0+x2/2

Shadow:

yeah so what is x, your variable on the left

Shadow:

what is that equation solving for

Shadow:

where did we get these equations from

lolokrat:

is this supposed to be an easy question or am i dumb

Shadow:

What formula did I post

Shadow:

What does it solve for

Shadow:

\[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\] this one

lolokrat:

midpoint

Shadow:

so what would x be

lolokrat:

the endpoint

lolokrat:

what

Shadow:

not x1 or x2, i'm talking about\[x = \frac{ x _{1} + x _{2} }{ 2 }\]

Shadow:

if x1 is the x coordinate of the first endpoint and x2 is the x coordinate of the second endpoint, then what is x?

Shadow:

What is that solving for?

lolokrat:

segment what i dont know this hurts my brain

Shadow:

the x coordinate of the midpoint

lolokrat:

the coordinates of the midpoint

Shadow:

yep

lolokrat:

r u serious

Shadow:

lol

Shadow:

you have that so I'm sure you know what to do now

Shadow:

solve for x2

Shadow:

my ride is here so I'll be on mobile later

lolokrat:

ok thanks for the help

Shadow:

You got it?

lolokrat:

i got (18,20) as my answer

Shadow:

Check your work again for y

lolokrat:

wow, i skpped a whole step. (18, 9)

Shadow:

You got it, good job

lolokrat:

thanks! can you check my others?

lolokrat:

endpoint: (13,12) midpoint: (7.5,5) answer: (2,-2)

Shadow:

Recheck x

lolokrat:

are you sure? what did i do wrong?

lolokrat:

7.5=13+x2/2 7.5*2=13+x2 15-13=x2 2=x2

Shadow:

Nvm you’re good.

lolokrat:

oh good

lolokrat:

2 more?

lolokrat:

endpoint: (-10,-2) midpoint: (-2.5, -9.5) answer: (-5, -21)

Shadow:

\[x = \frac{ x _{1} + x _{2} }{ 2 })\] \[-2.5 = \frac{ -10 + x _{2} }{ 2 }\] \[2 \times -2.5 = \frac{ -10 + x _{2} }{ 2 }\] \[-5 = -10 + x _{2} \] \[5 = x_{2}\]

Shadow:

-19 + 2 = -17 You made the same mistake. \[a - 1 = -2\] When you deal with situations like this when there is a negative on both sides and you are isolating the variable, you want to cancel out the constant by the variable, in this case it is -1. So you add 1 to both sides, not -1. If you add -1 to both sides you get: \[a -2 = -3\] That is what you were doing for the last step.

lolokrat:

ah, i always get confused with those negatives

Shadow:

Yeah the best way to think about it is as a number line.  If its a negative number, when you add 2 it becomes more positive, if you add -2 it becomes more negative. That's why subtraction can be rewritten like this: \[a - b = a + -b\]

lolokrat:

endpoint: (7, -7) midpoint: (17.5, -0.5) answer : (28,-8)

Shadow:

2 times -0.5 = -1 -1 + 7 = 6

Shadow:

Similar mistake as to earlier

lolokrat:

ah, sorry